OpenStudy (akonkel):

ANOTHER MEDAL AND FAN :) What is the least common multiple of 70, 60, and 50?

jagr2713 (jagr2713):

70 = 2 * 5 * 7 60 = 2 * 2 * 3 * 5 50 = 2 * 5 * 5 L.C.M. of 70, 60, and 50 = 2 * 2 * 3 * 5 * 5 * 7 = 2100

OpenStudy (akonkel):

@jagr2713 I GOT MORE MEDALS I WILL TAG U ;)